-20-3x+3x^2=0
I can't seem to find the right way to factor it!
: -20-3x+3x^2=0
I can't seem to find the right way to factor it!

Answer by Earlsdon(3748)

Perhaps that is because the trinomial -20-3x+3x^2=0 is not factorable...in other words, it's prime.
Let's rearrange the trinomial so that it's in standard form:
3x^2-3x-20 = 0
The factors, if there were any, would take the form:
(3x+m)(x+n) = 0 and you need to have m*n = -20 and 3n+m = -3
The factors of -20 are:
-20 = -1*20 or 1*(-20) and 3(-1)+20 = 17, 3(1)+(-20) = -17
-20 = -2*10 or 2*(-10) and 3(-2)+10 = 4, 3(2)+(-10) = -4
-20 = -4*5 or 4*(-5) and 3(-4)+5 = -7, 3(4)+(-5) = 7
So, as you can see, there is no way to get a -3 from combining the factors of -20.
The trinomial is prime!